# GPS-IDS: An Anomaly-based GPS Spoofing Attack Detection Framework for Autonomous Vehicles

### v1

@@ III-C5 Controller
III-D Attacker Model
@@
III-D1 Mathematical model of the spoofed GPS signal
@@
$y(t)=\textbf{Re}\Bigl{\{}\sum_{i=1}^{N}A_{i}D_{i}[t-\tau_{i}(t)]C_{i}[t-\tau_{i}(t)]e^{j\varphi}\Bigl{\}}$
$\varphi=\omega_{c}t-\phi_{i}(t)$
@@
$y_{spf}(t)=\textbf{Re}\Bigl{\{}\sum_{i=1}^{N_{spf}}A_{i_{spf}}\hat{D_{i}}[t-\tau_{i_{spf}}(t)]C_{i}[t-\tau_{i_{spf}}(t)]e^{j\hat{\varphi}}\Bigl{\}}$
$\hat{\varphi}=\omega_{c}t-\phi_{i_{spf}}(t)$
where $N_{spf}$ is the number of spoofed signals (typically $N_{spf}=N$). Each spoofed signal must have the same spreading code $C_{i}(t)$ as the corresponding true signal in order to deceive the receiver, and usually, it broadcasts its best estimate of the same data bit stream $\hat{D}_{i}(t)$. The spoofed amplitudes, code phases, and carrier phases are, respectively, $A_{i_{spf}}$, $\tau_{i_{spf}}(t)$, and $\phi_{i_{spf}}(t)$ for $i=1,...,N_{spf}$. These quantities are likely to differ from their true counterparts for reasons that are specific to the type of attack that is being mounted. During a spoofing attack, the total signal received, $y_{total}(t)$ at the victim vehicle’s GPS receiver antenna is:
@@
where $\nu(t)$ is the noise component, which can either be naturally generated or be contributed by the attacker.
III-D2 Mathematical model of the attack impacts
If the attack is successful, the GPS receiver of the vehicle will receive incorrect location information, and the current latitude and longitude will be replaced by spoofed latitude and longitude. The function $TargetYawFromGPS()$ in Algorithm 1 takes the spoofed latitude and longitude as inputs and calculates the spoofed yaw angle, $\psi_{t_{spf}}$. This means that the proportionality of equation 18 calculates a spoofed cross-track error, $e_{spf}(t)$. The PID controller takes $e_{spf}(t)$ as input and equation 19 and 20 become:
@@ III-D2 Mathematical model of the attack impacts
   Spoofed state $\textbf{x}_{spf}$, current_latitude, current_longitude
@@
1:  Attacker sends $y_{spf}$
@@ III-E Problem Statement
IV Experimental Evaluation
To validate the proposed GPS-IDS framework and demonstrate its effectiveness, field experiments have been conducted using an AV robotic testbed. The testbed has been designed and developed by following the Autonomous Vehicle Behavior Model presented in Fig. 4. This testbed has been utilized to perform GPS spoofing attack experiments and collect relevant data. The collected datasets have been used to conduct a series of experiments and validate the GPS-IDS framework.
IV-A Autonomous Vehicle Testbed (AVT)
The Autonomous Vehicle Testbed or AVT is a custom-built autonomous rover that can navigate through a predefined path using GPS guidance. It utilizes an array of state-measuring sensors, including accelerometer, magnetometer, barometer, gyroscope, and digital compass for accurate state estimation. In accordance with the Autonomous Vehicle Behavior Model, the AVT employs a PID controller and a multi-sensor fusion-based EKF failsafe. The EKF failsafe is triggered when the EKF variances associated with any two state-measuring sensor readings exceed a predefined EKF threshold value for 1 second. The AVT uses Ardupilot [64], an open-source software and hardware platform designed for building custom unmanned ground and aerial robotic vehicles. An Ardupilot-based autopilot generates the velocity and steering angle commands from the GPS and feedback from the state-measuring sensors of the AVT. These commands are then passed to the PID controller for execution. To program the AVT and define a path to follow using GPS, a Ground Control Station software supported by Ardupilot was utilized. The Ground Control Station computer receives all vehicular data via the telemetry modules in real time and stores them in the local memory of the computer. These vehicular data are referred to as “Dataflash logs”, and collected at Ardupilot’s default data update rate of 1 Hz.
@@ IV-A Autonomous Vehicle Testbed (AVT)
IV-B GPS Spoofing Attack on the AVT
To perform the GPS spoofing attack on the AVT system, a real-life spoofing attack scenario was generated based on the Attacker Model outlined in Section III-D of this paper. The attack experiment utilized a single spoofed signal, mathematically denoted as $N_{spf}$ = 1. Consistent with the standard L1 GPS signal, the spoofed signal employed a carrier wave frequency of 1575.42 MHz. The navigation message contained information specific to a single spoofed location, which was modulated onto the carrier wave using the BPSK modulation technique. As outlined in Algorithm 2, the spoofing attack affected the AVT control system and compromised its normal state.
IV-C Data Collection
The datasets were collected in two distinct locations separated by approximately 4 miles: The University of Arizona campus, and the Alvernon Park— both located in Tucson, Arizona, USA. During data collection, the vehicle was consistently operated in autonomous mode guided by GPS. As a safety precaution, a wireless remote controller was employed for manual control of the vehicle to mitigate potential hazards.
@@ IV-C Data Collection
IV-D Autonomous Vehicle GPS Dataset (AV-GPS-Dataset)
The AV-GPS-Dataset is composed of three subsets, each with varied entries, namely AV-GPS-Dataset 1, AV-GPS-Dataset 2, and AV-GPS-Dataset 3. Each subset contains two classes of data: Normal and Attack data, and they are extracted as Comma Separated Value (CSV) files from the Ground Control Station computer. The AV-GPS-Dataset mainly consists of temporal features of the AVT. The features and labels are consistent across all datasets, ensuring uniformity in the categorization of the data.
@@ IV-D Autonomous Vehicle GPS Dataset (AV-GPS-Dataset)
IV-E Experimental Analysis
@@
IV-E1 Experiment 1: State Estimation of the AVT using the Dynamic Bicycle Model
This experiment establishes a connection between the Autonomous Vehicle Behavior Model and the actual AVT. In this experiment, we estimated the states of the AVT using the dynamic bicycle model presented in section III-C1. The estimated results are then compared with the actual vehicle states obtained from 46,787 instances of normal data from AV-GPS-Dataset 1. Since a small-scale testbed is used to represent the real vehicle, it is assumed that the distances of the front and the rear wheel axle from the center of mass are equal. It is also assumed that the cornering stiffness coefficients for the front and rear wheels are equal to 1. Table I presents the physical parameters of the AVT that were used to identify the system matrix “a” and the input matrix “b”. Substituting these values in Equation 15, the final dynamic state equation obtained is as follows:
@@ IV-E1 Experiment 1: State Estimation of the AVT using the Dynamic Bicycle Model
The estimated results of the lateral dynamics of the AVT compared with the actual values are shown in Fig. 6. From Fig. 6, it can be observed that the estimated states using the dynamic bicycle model follow a similar distribution as the actual values of the yaw rates and lateral velocities. Therefore, the derived dynamic bicycle model can represent the lateral dynamics of the AVT well. This points out the necessity of adapting the dynamic bicycle model to capture the behavior of the autonomous vehicle and collect data on the parameters identified by the vehicle behavior model. Since the presented vehicle model is capable of estimating the distributions of the next states of the AVT correctly, the state space representation accurately models the testbed. The differences in these values can be accepted based on the assumption that the field experiments added some noise that is not considered in the state equation.
@@
Figure 7: AVT pose measurements and controller input/ output without imposing GPS spoofing attack
(e) Position measurement
(f) Orientation measurement
(g) Controller input
(h) Controlled speed
@@
IV-E2 Experiment 2: Attack Impact on Pose Measurements and Controller Input/ Output
This experiment presents a comprehensive analysis of the behavior of the AVT under normal operating conditions and under the influence of GPS spoofing attack. We initially depicted the AVT’s normal behavior in terms of pose measurements and controller input/ output by plotting one round of normal data, as shown in Fig. 8. Fig. 7(a) and Fig. 7(b) illustrate the AVT’s position and orientation, while Fig. 7(c) and Fig. 7(d) illustrate the cross-track error and velocity, respectively, under normal operating conditions. From Fig. 8, it is observed that the AVT follows a circular trajectory with a velocity ranging from 0 m/s to 2.5 m/s. During its path traversal, it effectively corrects any cross-track errors and adjusts its yaw angles accordingly.
To facilitate a clear comparison of the attack impact, we then plotted one round of normal data with the corresponding attack data pertaining to the pose measurements and controller input/ output of the AVT, as presented in Fig. 8. In Fig. 8, it is apparent that the AVT’s position is significantly affected by the spoofed GPS signals. It deviates from its regular circular trajectory and moves toward a different direction along the X coordinate (Fig. 7(e)) at an irregular velocity spiking to as high as 80 m/s, and then decelerating to 0 m/s (Fig. 7(h)). In Fig. 7(g), we can observe an abnormal rise in cross-track errors exceeding 740 meters, then dropping to around 0 meters after 35 seconds and continuing to be the same. The AVT was unable to correct this error and adjust its heading, causing the yaw angle to remain unchanged (Fig. 7(f)). During this time, the AVT had lost all the satellite locks and remained stationary due to the variance in EKF estimates. Nonetheless, on the Ground Control Station map, the AVT was observed virtually progressing toward the location indicated by the spoofed signals, which is shown in Fig. 9.
TABLE II: Performance of the Machine Learning Classification Models on AV-GPS-Datasets
AV-GPS-Datasets Metrics RF XGB SVC MLP Adaboost GB DT
@@ IV-E2 Experiment 2: Attack Impact on Pose Measurements and Controller Input/ Output
IV-E3 Experiment 3: Performance Analysis of the Machine Learning Models with Different Training Sets
This experiment evaluated the GPS-IDS approach on the AV-GPS-Dataset using different machine learning models. To benchmark various machine learning techniques, including ensemble methods, neural networks, and tree-based algorithms, seven models were chosen for detection: Random Forest (RF), XGBoost (XGB), Support Vector Machine Classifier (SVC), Multi-Layer Perceptron (MLP), AdaBoost, Gradient Boosting (GB), and Decision Tree (DT).
Features associated with equations 15 (Vehicle Model block), 16 (Localization block), 17 (State Estimation block), 18 (Motion Planning block), and 20 (PID Controller block) from the Autonomous Vehicle Behavior Model (Fig. 4) are extracted and utilized in the machine learning models. To maintain the proportion of samples for each class consistently across both training and testing phases, stratified 5-fold sampling was employed to partition the data into training and testing sets. The training approach is based on supervised learning, where 4 folds (80%) were utilized for training, and the remaining fold (20%) was utilized for testing. Three cases were considered to train the models, Case I: Train on 80% of AV-GPS-Dataset 1, and test on 20% of AV-GPS-Dataset 1, AV-GPS-Dataset 2 and AV-GPS-Dataset 3; Case II: Train on 80% of AV-GPS-Dataset 2, and test on AV-GPS-Dataset 1, 20% of AV-GPS-Dataset 2 and AV-GPS-Dataset 3; Case III: Train on 80% of AV-GPS-Dataset 3, and test on AV-GPS-Dataset 1, AV-GPS-Dataset 2 and 20% of AV-GPS-Dataset 3. For all three cases, the performance of the machine learning models is measured in terms of Accuracy, Precision, Recall, and F1-score. For each model, a combination of hyperparameters was utilized and fine-tuned to find the optimal performance. The performance of machine learning algorithms on different datasets is presented in Table II. The results indicate that the performance varies across different datasets as the training set changes. It can be observed that Case I exhibits high accuracy and F1 scores for most models, while the F1 scores drop for most models in Case III. This drop can be attributed to the smaller size of the training set in Case III, which contains a limited number of entries representing only the transition of the attack. In this case, the test sets are significantly larger than the training set (AV-GPS-Dataset 1 is over 110 times larger, and AV-GPS-Dataset 2 is over 12 times larger), which declines the model performances. Overall, MLP consistently performs with F1 scores above 90% in all three cases, achieving the highest average F1 score of 94.4%. Following closely, RF and XGB demonstrated similar and second-best performances, achieving average F1 scores of 91.9% and 91.5%, respectively, in the three considered cases. The DT classifier secured the third-best performance, attaining an average F1 score of 89.1% across the three cases.
IV-E4 Experiment 4: Tuning of the Detection Threshold $\mathbb{T}$

### v2

@@ III-B Anomaly Behavior Analysis (ABA) Module
IV Attacker Model and Problem Statement
@@
IV-A Mathematical model of the spoofed GPS signal
@@
$y(t)=\textbf{Re}\Bigl{\{}\sum_{i=1}^{N}A_{i}D_{i}[t-\tau_{i}(t)]C_{i}[t-\tau_{i}(t)]e^{j\varphi}\Bigl{\}}$ (29)
$\varphi=\omega_{c}t-\phi_{i}(t)$ (30)
@@
$y_{spf}(t)=\textbf{Re}\Bigl{\{}\sum_{i=1}^{N_{spf}}A_{i_{spf}}\hat{D_{i}}[t-\tau_{i_{spf}}(t)]C_{i}[t-\tau_{i_{spf}}(t)]e^{j\hat{\varphi}}\Bigl{\}}$ (31)
$\hat{\varphi}=\omega_{c}t-\phi_{i_{spf}}(t)$ (32)
where $N_{spf}$ is the number of spoofed signals (typically $N_{spf}=N$). Each spoofed signal must have the same spreading code $C_{i}(t)$ as the corresponding true signal in order to deceive the receiver, and usually, it broadcasts its best estimate of the same data bit stream $\hat{D}_{i}(t)$. The spoofed amplitudes, code phases, and carrier phases are, respectively, $A_{i_{spf}}$, $\tau_{i_{spf}}(t)$, and $\phi_{i_{spf}}(t)$ for $i=1,...,N_{spf}$. According to the mathematical model, when a spoofing attack is successful, the victim vehicle’s GPS receiver antenna receives $y(t)+y_{spf}(t)$.
This attack is still detectable by several approaches discussed in II section. To enhance the stealthiness of our GPS spoofing attack, realistic noise was added to the generated spoofed signals by sampling a Gaussian distribution with a large variance, as detailed by Wang et al. [67]. The addition of Gaussian noise aims to mimic the natural variations and imperfections present in genuine GPS signals, making it more challenging for IDSs to identify the spoofing attack.
Finally, the total signal received, $y_{total}(t)$ at the victim vehicle’s GPS receiver antenna is:
@@
where $\nu(t)$ is the Gaussian noise component added to the spoofing attack to make it stealthy.
IV-B Mathematical model of the attack impacts on the Vehicle
If the attack is successful, the GPS receiver of the vehicle will receive incorrect location information, and the current latitude and longitude will be replaced by spoofed latitude and longitude. The function $TargetYawFromGPS()$ in Algorithm 1 takes the spoofed latitude and longitude as inputs and calculates the spoofed yaw angle, $\psi_{t_{spf}}$. This means that the proportionality of equation 25 calculates a spoofed cross-track error, $e_{spf}(t)$. The PID controller takes $e_{spf}(t)$ as input and equation 26 and 27 become:
@@ IV-B Mathematical model of the attack impacts on the Vehicle
   Spoofed state $\textbf{x}_{spf}$, Gaussian noise component $\nu(t)$,
   current_latitude, current_longitude
@@
1:  Attacker sends $y_{spf}+\nu\Rightarrow\text{stealthy attack}$
@@ IV-C Problem Statement
V Experimental Evaluation
To validate the proposed GPS-IDS framework and demonstrate its effectiveness, field experiments were conducted using an AV robotic testbed. The testbed was designed and developed by following the Autonomous Vehicle Behavior Model presented in Fig. 4. This testbed was utilized to perform GPS spoofing attack experiments and collect relevant real-world data. In addition, simulated data with urban environment scenarios were generated using CARLA simulator. All the collected datasets have been used to conduct a series of experiments and validate the GPS-IDS framework.
V-A Field Experiments and Data Collection
V-A1 Autonomous Vehicle Testbed (AVT)
The Autonomous Vehicle Testbed or AVT is a custom-built autonomous rover that can navigate through a predefined path using GPS guidance. It utilizes an array of state-measuring sensors, including an accelerometer, magnetometer, barometer, gyroscope, and digital compass for accurate state estimation. Following the Autonomous Vehicle Behavior Model, the AVT employs a PID controller and a GPS/INS fusion-based EKF failsafe. The EKF failsafe is triggered when the EKF variances associated with any two state-measuring sensor readings exceed a predefined EKF threshold value for 1 second. The AVT uses Ardupilot [64], an open-source software and hardware platform designed for building custom unmanned ground and aerial robotic vehicles. An Ardupilot-based autopilot generates the velocity and steering angle commands from the GPS and feedback from the state-measuring sensors of the AVT. These commands are then passed to the PID controller for execution. To program the AVT and define a path to follow using GPS, a Ground Control Station software supported by Ardupilot was utilized. The Ground Control Station computer receives all vehicular data via the telemetry modules in real time and stores them in the local memory of the computer. These vehicular data are referred to as “Dataflash logs”, and collected at Ardupilot’s default data update rate of 1 Hz.
@@ V-A1 Autonomous Vehicle Testbed (AVT)
V-A2 Practical GPS Spoofing Attack on the AVT
To perform the GPS spoofing attack on the AVT, a real-life spoofing attack scenario was generated based on the Attacker Model outlined in Section IV of this paper. The attack experiment utilized a single spoofed signal, mathematically denoted as $N_{spf}$ = 1. Consistent with the standard L1 GPS signal, the spoofed signal employed a carrier wave frequency of 1575.42 MHz. The navigation message contained information specific to a spoofed location, which was modulated onto the carrier wave using the BPSK modulation technique. The GPS satellite constellation was specified via a GPS broadcast ephemeris file (a collection of the individual site navigation files). The sampling frequency was set to 2600 MHz and a Gaussian noise of 20 MHz centered around 1575.42 MHz was generated for a stealthy attack. The spoofed signal was generated using the GPS-SDR-SIM tool [76] and transmitted at a rate of 1 Hz.
To ensure ethical compliance and prevent interference with other GPS receivers or critical navigation systems, the transmission range and strength of the spoofed signal were strictly controlled and contained — limiting it to less than 3 meters by controlling the transmit power. A test receiver was placed at various distances to measure the range. The transmit gain was set very low ($\approx$ -10 dB) and then gradually incremented while monitoring the spoofed signal’s range. The gain setting at which the signal was detectable at 3 meters and dropped off significantly beyond that range was noted ($\approx$ 0 dB) and utilized in the experiments. The spoofing experiments were conducted in open, wide fields with no other vehicles or critical systems within range. As outlined in Algorithm 2, the spoofing attack affected the AVT control system and compromised its normal state.
V-A3 Real Data Collection
The real-world datasets were collected in two distinct locations separated by approximately 4 miles: The University of Arizona campus, and the Alvernon Park— both located in Tucson, Arizona, USA. During data collection, the vehicle was consistently operated in autonomous mode guided by GPS. As a safety precaution, a wireless remote controller was employed for manual control of the vehicle to mitigate potential hazards.
@@ V-A3 Real Data Collection
V-B Simulation Experiments
V-B1 CARLA Simulator
CARLA is an open-source urban driving simulator designed to facilitate the development, training, and validation of autonomous driving systems. CARLA provides a highly detailed, realistic environment that mimics urban settings, complete with dynamic weather conditions, diverse vehicle and pedestrian traffic, and a wide array of sensory inputs. We utilized CARLA to run simulations in an urban environment where various vehicle sensor data was collected, including GPS signals, accelerometer readings, and gyroscope data.
V-B2 Simulation Environment
The simulations were conducted in CARLA Town01, which is a small, urban-like area featuring a mix of streets, a few T-junctions, and a variety of buildings. The town is surrounded by coniferous trees and contains several small bridges that span a river, dividing the town into two halves. We ran 3 simulations, each lasting 5 minutes, to capture various randomizations of vehicle start locations. The simulations were performed in an urban town setting with clear weather conditions. In each simulation, 100 vehicles were spawned at different locations. Among these, 10 vehicles were equipped with GPS receivers, accelerometers, and gyroscopes. All vehicles in the simulations were controlled by an autopilot system to generate subsequent locations. The sensor data from the 10 equipped vehicles were monitored and logged at 0.1-second intervals.
V-B3 GPS Spoofing Attack Simulation
To simulate the GPS spoofing attack, we modified the GPS spoofed entries according to the equation: $X^{\prime}\sim\mathcal{N}(X+\mu,\,\sigma^{2})\,.$ where $X^{\prime}$ is the GPS reading (latitude and longitude) after adding the spoofing effect, $X$ is the GPS reading before adding the spoofing effect, $\mu$ the mean of the GPS readings for each vehicle and $\sigma$ is the standard deviation of the GPS readings for each vehicle. This GPS spoofing method was used in [66]. We generated a GPS spoofed entry for each original entry.
V-C Autonomous Vehicle GPS Datasets (AV-GPS-Datasets)
The AV-GPS-Datasets have two parts with a total of 6 subsets, namely AV-GPS-Dataset (Real Data), with subsets 1—3, and AV-GPS-Dataset (Simulation-generated Data), with subsets 4—6. Each subset contains two classes of data: Normal and Attack data, and they are extracted as Comma Separated Value (CSV) files. The AV-GPS-Datasets mainly consist of temporal features of AVs collected using the AVT and CARLA simulated environment.
V-C1 AV-GPS-Dataset 1—3 (Real Data)
@@
V-C2 AV-GPS-Dataset 4—6 (Simulation-generated Data)
To prepare AV-GPS-Dataset 4—6 from CARLA simulations, the vehicle sensor data was first collected in normal conditions without imposing GPS spoofing attacks. Data from the 14 features highlighted in the Autonomous Vehicle Behavior Model was collected. Three simulations were conducted, resulting in three distinct datasets: AV-GPS-Dataset 4 with 29,735 entries of normal data, AV-GPS-Dataset 5 with 22,120 entries of normal data, and AV-GPS-Dataset 6 with 21,315 entries of normal data. Each dataset features different random starting positions for each vehicle. Traffic density varied across the datasets, with AV-GPS-Dataset 4 having heavy traffic density, AV-GPS-Dataset 5 having moderate traffic density, and AV-GPS-Dataset 6 having sparse traffic density. We then introduced GPS spoofing attacks by adding corresponding spoofed data entries as explained in V-B3 and reran the experiments to collect the attack data, effectively doubling the number of data points in each dataset. As a result, AV-GPS-Dataset 4 contains a total of 59,470 data points, AV-GPS-Dataset 5 contains 44,240 data points, and AV-GPS-Dataset 6 contains 42,630 data points.
V-D Experimental Analysis
@@
V-D1 Experiment 1- State Estimation of the AVT using the Autonomous Vehicle Behavior Model
This experiment establishes a connection between the Autonomous Vehicle Behavior Model and the actual AVT. In this experiment, we estimated the states of the AVT using the dynamic bicycle model presented in section III-A3. The estimated results are then compared with the actual vehicle states obtained from 46,787 instances of normal data from AV-GPS-Dataset 1. Since a small-scale testbed is used to represent the real vehicle, it is assumed that the distances of the front and the rear wheel axle from the center of mass are equal. It is also assumed that the cornering stiffness coefficients for the front and rear wheels are equal to 1. Table I presents the physical parameters of the AVT that were used to identify the system matrix “a” and the input matrix “b”. Substituting these values in Equation 13, the final dynamic state equation obtained is as follows:
@@ V-D1 Experiment 1- State Estimation of the AVT using the Autonomous Vehicle Behavior Model
The estimated results of the lateral dynamics of the AVT compared with the actual values are shown in Fig. 6. From Fig. 6, it can be observed that the estimated states using the dynamic bicycle model follow a similar distribution as the actual values of the yaw rates and lateral velocities. Therefore, the derived dynamic bicycle model can represent the lateral dynamics of the AVT well. This points out the necessity of adapting the dynamic bicycle model to capture the behavior of AVs and collect data on the parameters identified by the vehicle behavior model. Since the presented vehicle model is capable of estimating the distributions of the next states of the AVT correctly, the state space representation accurately models the testbed. The differences in these values can be accepted based on the assumption that the field experiments added some noise that is not considered in the state equation.
@@
V-D2 Experiment 2- Attack Impact on Pose Measurements and Controller Input/ Output
This experiment presents a comprehensive analysis of the behavior of the AVT under normal operating conditions and under the influence of real GPS spoofing attack. To facilitate a clear comparison of the attack impact, we plotted one round of normal data with the corresponding attack data from AV-GPS-Dataset 1 pertaining to the pose measurements and controller input/ output of the AVT, as presented in Fig. 7. We can observe that the AVT follows a circular trajectory (7(a)) with a velocity ranging from 0 m/s to 2.5 m/s. During its path traversal, it effectively corrects cross-track errors and adjusts its yaw angles accordingly (7(d)).
It is apparent that the AVT’s position is significantly affected by the spoofed GPS signals. It deviates from its regular circular trajectory and moves toward a different direction along the X coordinate (Fig. 7(a)) at an irregular velocity spiking to as high as 80 m/s, and then decelerating to 0 m/s (Fig. 7(d)). In Fig. 7(c), we can observe an abnormal rise in cross-track errors exceeding 740 meters, then dropping to around 0 meters after 35 seconds and continuing to be the same. The AVT was unable to correct this error and adjust its heading, causing the yaw angle to remain unchanged (Fig. 7(b)). During this time, the AVT had lost all the satellite locks and remained stationary due to the variance in EKF estimates. Nonetheless, on the Ground Control Station map, the AVT was observed virtually progressing toward the location indicated by the spoofed signals, which is shown in Fig. 8.
TABLE II: Performance of the Machine Learning Classification Models on AV-GPS-Datasets 1—3 (Real Data)
AV-GPS-Datasets 1—3 Metrics RF XGB SVC MLP Adaboost GB DT
@@ V-D2 Experiment 2- Attack Impact on Pose Measurements and Controller Input/ Output
V-D3 Experiment 3- Performance Analysis of the Machine Learning Models with Different Training Sets from Real Data
This experiment evaluated the GPS-IDS approach on the AV-GPS-Dataset 1—3 using different machine learning models. To benchmark various machine learning techniques, including ensemble methods, neural networks, and tree-based algorithms, seven models were chosen for detection: Random Forest (RF), XGBoost (XGB), Support Vector Machine Classifier (SVC), Multi-Layer Perceptron (MLP), AdaBoost, Gradient Boosting (GB), and Decision Tree (DT). The 14 features associated with the physics-based component of the Autonomous Vehicle Behavior Model (Fig. 4) are utilized in the machine learning analysis. To maintain the proportion of samples for each class consistently across both training and testing phases, stratified 5-fold sampling was employed to partition the data into training and testing sets. The training approach is based on supervised learning, where 4 folds (80%) were utilized for training, and the remaining fold (20%) was utilized for testing. Three cases were considered to train the models, Case I: Train on 80% of AV-GPS-Dataset 1, and test on 20% of AV-GPS-Dataset 1, AV-GPS-Dataset 2 and AV-GPS-Dataset 3; Case II: Train on 80% of AV-GPS-Dataset 2, and test on AV-GPS-Dataset 1, 20% of AV-GPS-Dataset 2 and AV-GPS-Dataset 3; Case III: Train on 80% of AV-GPS-Dataset 3, and test on AV-GPS-Dataset 1, AV-GPS-Dataset 2 and 20% of AV-GPS-Dataset 3. For all three cases, the performance of the machine learning models is measured in terms of Accuracy, Precision, Recall, and F1-score. For each model, a combination of hyperparameters was utilized and fine-tuned to find the optimal performance. The performance of machine learning algorithms on different datasets is presented in Table II. The results indicate that the performance varies across different datasets as the training set changes. It can be observed that Case I exhibits high accuracy and F1 scores for most models, while the F1 scores drop for most models in Case III. This drop can be attributed to the smaller size of the training set in Case III, which contains a limited number of entries representing only the transition of the attack. In this case, the test sets are significantly larger than the training set (AV-GPS-Dataset 1 is over 110 times larger, and AV-GPS-Dataset 2 is over 12 times larger), which declines the model performances. Overall, MLP consistently performs with F1 scores above 90% in all three cases, achieving the highest average F1 score of 94.4%. Following closely, RF and XGB demonstrated similar and second-best performances, achieving average F1 scores of 91.9% and 91.5%, respectively, in the three considered cases. The DT classifier secured the third-best performance, attaining an average F1 score of 89.1% across the three cases.
V-D4 Experiment 4- Tuning of the Detection Threshold $\mathbb{T}$